# Copula‐Stein Discrepancy: A Generator‐Based Stein Operator for Archimedean Dependence

### v1

@@ 5.2 A practical test
Proof. From the witness function definition in Proposition 5.1:
@@
5.3 Application to Archimedean Copulas
Proposition 5.3 (Archimedean Score Function)
Let $C$ be an Archimedean copula on $(0,1)^{d}$ with generator $\varphi:(0,1]\to[0,\infty)$ strictly decreasing and twice differentiable (Assumptions 1), and let $\psi=\varphi^{-1}$. For $u\in(0,1)^{d}$, set
$t\;=\;\sum_{k=1}^{d}\varphi(u_{k}).$
Then the coordinate-wise score of the copula density $c$ is
$\frac{\partial}{\partial u_{j}}\log c(u)\;=\;\varphi^{\prime}(u_{j})\,\frac{\psi^{(d+1)}(t)}{\psi^{(d)}(t)}\;+\;\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}\;,\qquad j=1,\dots,d.$
Proof. The standard density form is
$c(u)\;=\;\psi^{(d)}(t)\;\prod_{k=1}^{d}\!\big(-\varphi^{\prime}(u_{k})\big),\qquad t=\sum_{k=1}^{d}\varphi(u_{k}).$
@@ 5.3 Application to Archimedean Copulas
$\frac{\partial}{\partial u_{j}}\log c(u)=\Big(\partial_{t}\log\psi^{(d)}(t)\Big)\,\frac{\partial t}{\partial u_{j}}\;+\;\frac{\partial}{\partial u_{j}}\log\!\big(-\varphi^{\prime}(u_{j})\big).$
@@
Since $\partial t/\partial u_{j}=\varphi^{\prime}(u_{j})$ and $\partial_{t}\log\psi^{(d)}(t)=\psi^{(d+1)}(t)/\psi^{(d)}(t)$ while $\frac{\partial}{\partial u_{j}}\log\!\big(-\varphi^{\prime}(u_{j})\big)=\varphi^{\prime\prime}(u_{j})/\varphi^{\prime}(u_{j})$, the claim follows.   We used $\frac{d}{du}\log(\varphi^{\prime}(u))=\varphi^{\prime\prime}(u)/\varphi^{\prime}(u)$ since $\varphi^{\prime}(u)<0$ for Archimedean generators.
Remark 5.2
$\frac{\partial}{\partial u_{j}}\log c(u)=\Big[\varphi^{\prime}(w)\,\partial_{t}\log\psi^{(d)}(t)\Big]\frac{\varphi^{\prime}(u_{j})}{\varphi^{\prime}(w)}+\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}=\varphi^{\prime}(u_{j})\,\partial_{t}\log\psi^{(d)}(t)+\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}.$
This analytic score makes the CSD kernel highly efficient. We can further show that for common choices of base kernels, the cost per evaluation is linear in the dimension.
Proposition 5.4 (Computational Complexity)
For separable kernels $k(u,v)=\prod_{j=1}^{d}k_{1}(u_{j},v_{j})$, each evaluation of the Copula-Stein kernel $k_{C_{0}}(u,v)$ requires $O(d)$ operations.
Proof. For a separable kernel, the gradients and Hessian have special structure:
$\frac{\partial k(u,v)}{\partial u_{j}}$ $=k_{1}^{\prime}(u_{j},v_{j})\prod_{i\neq j}k_{1}(u_{i},v_{i}),$ (25)
$\frac{\partial k(u,v)}{\partial v_{j}}$ $=k_{1}^{\prime}(u_{j},v_{j})\prod_{i\neq j}k_{1}(u_{i},v_{i}),$ (26)
$\frac{\partial^{2}k(u,v)}{\partial u_{j}\partial v_{j}}$ $=k_{1}^{\prime\prime}(u_{j},v_{j})\prod_{i\neq j}k_{1}(u_{i},v_{i}).$ (27)
The computational cost breakdown:
  1. 1.
    Computing $k(u,v)=\prod_{j=1}^{d}k_{1}(u_{j},v_{j})$: $O(d)$ operations
  2. 2.
    Computing score functions $s_{j}(u)$ for all $j$: $O(d)$ operations (using Proposition 3.4)
  3. 3.
    Computing gradient terms $\nabla_{u}k(u,v)$ and $\nabla_{v}k(u,v)$: $O(d)$ operations each
  4. 4.
    Computing trace term $\text{tr}(\nabla_{u}\nabla_{v}^{T}k(u,v))$: $O(d)$ operations
  5. 5.
    Final combination in kernel formula: $O(d)$ operations
Total: $O(d)$ operations per kernel evaluation.
6.  Theoretical Properties
6.1 Metrization and Consistency
Theorem 2 (Detectability)
Under Assumptions 1 (A1–A4), $\mathrm{CSD}(Q,C_{0})=0$ if and only if $Q=C_{0}$.
Lemma 2 (Stein solution in the RKHS)
Under A1A3, for any bounded–Lipschitz $h$ with $\mathbb{E}_{C_{0}}[h]=0$, there exists $g_{h}\in\mathcal{H}^{d}$ such that $\mathcal{T}_{C_{0}}g_{h}=h$ and $\|g_{h}\|_{\mathcal{H}^{d}}\leq C\|h\|_{\mathrm{BL}}$.
Proof. Construct $g_{h,1}$ by solving $\partial_{u_{1}}g_{h,1}+s_{1}g_{h,1}=h-\sum_{j\geq 2}(\partial_{u_{j}}g_{h,j}+s_{j}g_{h,j})$ with zero trace at $u_{1}\in\{0,1\}$; the growth bound on $s_{1}$ (A1) and Lipschitz $h$ yield a uniform bound. Approximate $g_{h,1}$ by elements of $\mathcal{H}$ using A3 and pass limits through $\mathcal{T}_{C_{0}}$ by dominated convergence.
Theorem 3 (Metrization of weak convergence)
Let $(Q_{n})$ be probability laws on $(0,1)^{d}$. Under A1A3,
$\mathrm{CSD}(Q_{n},C_{0})\to 0\quad\Longleftrightarrow\quad Q_{n}\Rightarrow C_{0}.$
Proof. The forward direction follows from boundedness and continuity of $\mathcal{T}_{C_{0}}g$ when $g\in\mathcal{H}^{d}$ and $\|g\|_{\mathcal{H}^{d}}\leq 1$, together with the Portmanteau theorem and the Stein identity under $C_{0}$ in A2. We prove the reverse direction.
Fix any bounded Lipschitz $h:(0,1)^{d}\to\mathbb{R}$. Write $\bar{h}\equiv h-\mathbb{E}_{C_{0}}[h]$. We claim there exists $g_{h}\in\mathcal{H}^{d}$ such that
$\mathcal{T}_{C_{0}}g_{h}=\bar{h}\qquad\text{and}\qquad\|g_{h}\|_{\mathcal{H}^{d}}\leq C\,\|h\|_{\mathrm{BL}},$ (28)
for a constant $C$ depending only on $c_{0}$ and $\mathcal{H}$. This is a Stein equation for the operator $\mathcal{T}_{C_{0}}$.
We construct $g_{h}$ by solving a first order linear equation componentwise along the $u_{1}$ coordinate with zero boundary data. Let $s_{1}=\partial_{u_{1}}\log c_{0}$. Define, for $u\in(0,1)^{d}$,
$g_{h,1}(u)\equiv\int_{0}^{u_{1}}\bar{h}(t,u_{2:d})\,\exp\!\Big\{\int_{t}^{u_{1}}s_{1}(r,u_{2:d})\,dr\Big\}\,dt,\qquad g_{h,j}\equiv 0\ \text{for }j\geq 2.$
A direct differentiation shows $\partial_{u_{1}}g_{h,1}+s_{1}g_{h,1}=\bar{h}$, hence $\mathcal{T}_{C_{0}}g_{h}=\bar{h}$. The boundary value $g_{h,1}(0,u_{2:d})=0$ and $g_{h,1}(1,u_{2:d})$ is finite due to A1 and the boundedness of $h$.
We bound $\|g_{h}\|_{\mathcal{H}^{d}}$. Since $\mathcal{H}$ is dense in $C^{1}$ on compact subsets by A3, approximate the scalar map $u_{1}\mapsto g_{h,1}(u_{1},u_{2:d})$ uniformly by $\mathcal{H}$ functions for each fixed $u_{2:d}$ and pass limits under the integral sign using dominated convergence, valid under A1. Standard RKHS interpolation bounds then give
$\|g_{h,1}\|_{\mathcal{H}}\leq C_{1}\|\bar{h}\|_{\infty}+C_{2}\mathrm{Lip}(h)\leq C\,\|h\|_{\mathrm{BL}},$
where $C$ depends on $\sup|s_{1}|$ on compact subsets and on the kernel constants. Hence $g_{h}\in\mathcal{H}^{d}$ and (28) holds.
By (28) and the Stein identity,
$\mathbb{E}_{Q_{n}}[h]-\mathbb{E}_{C_{0}}[h]=\mathbb{E}_{Q_{n}}[\bar{h}]=\mathbb{E}_{Q_{n}}[\mathcal{T}_{C_{0}}g_{h}]\leq\mathrm{CSD}(Q_{n},C_{0})\,\|g_{h}\|_{\mathcal{H}^{d}}\leq C\,\|h\|_{\mathrm{BL}}\,\mathrm{CSD}(Q_{n},C_{0}).$
Since $\mathrm{CSD}(Q_{n},C_{0})\to 0$, we get $\mathbb{E}_{Q_{n}}[h]\to\mathbb{E}_{C_{0}}[h]$ for all bounded Lipschitz $h$. This is equivalent to $Q_{n}\Rightarrow C_{0}$.
6.2 Finite‐Sample Convergence and Optimality (Detailed Proof)
Theorem 4 (Finite‐Sample Convergence Rate)
Under Assumptions 1, one has
and hence $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$.
Proof. We write
@@ 6.2 Finite‐Sample Convergence and Optimality (Detailed Proof)
together with Lemma 1, imply $\mathbb{E}[k(U_{i},U_{j})]=0$. Hence
@@
as claimed.
Corollary 6.1 (Optimal High‐Probability Convergence Rate)
Under the unified assumptions, the CSD estimator satisfies
$\widehat{\mathrm{CSD}}_{n}=O_{P}\bigl(n^{-1/2}\bigr).$ (36)
Proof. Decompose $\widehat{\mathrm{CSD}}_{n}^{2}=D_{\text{diag}}+D_{\text{off}}$ with

### v2

@@ 5.2 A practical test
$k_{C_{0}}(u,v)$ $=s(u)^{T}s(v)k(u,v)$ (22)
$+s(u)^{T}\nabla_{v}k(u,v)+s(v)^{T}\nabla_{u}k(u,v)$ (23)
$+\text{tr}(\nabla_{u}\nabla_{v}^{T}k(u,v)).$ (24)
Full proof is provided in Appendix A.4.
5.3 Application to Archimedean Copulas
Proposition 5.7 (Archimedean Score Function).
Let $C$ be an Archimedean copula on $(0,1)^{d}$ with generator $\varphi:(0,1]\to[0,\infty)$ strictly decreasing and twice differentiable (Assumptions 4.2), and let $\psi=\varphi^{-1}$. For $u\in(0,1)^{d}$, set
$t\;=\;\sum_{k=1}^{d}\varphi(u_{k}).$ (25)
Then the coordinate-wise score of the copula density $c$ is
$\;\frac{\partial}{\partial u_{j}}\log c(u)\;=\;\varphi^{\prime}(u_{j})\,\frac{\psi^{(d+1)}(t)}{\psi^{(d)}(t)}\;+\;\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}\;,$ (26)
$j=1,\dots,d.$ (27)
Proof is provided in Appendix A.5.
We used $\frac{d}{du}\log(\varphi^{\prime}(u))=\varphi^{\prime\prime}(u)/\varphi^{\prime}(u)$ since $\varphi^{\prime}(u)<0$ for Archimedean generators.
Remark 5.8.
$\frac{\partial}{\partial u_{j}}\log c(u)$ $=\Big[\varphi^{\prime}(w)\,\partial_{t}\log\psi^{(d)}(t)\Big]\frac{\varphi^{\prime}(u_{j})}{\varphi^{\prime}(w)}+\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}$ (28)
$=\varphi^{\prime}(u_{j})\,\partial_{t}\log\psi^{(d)}(t)+\frac{\varphi^{\prime\prime}(u_{j})}{\varphi^{\prime}(u_{j})}.$ (29)
6 Theoretical Properties
6.1 Metrization and Consistency
Theorem 6.1 (Detectability).
Under Assumptions 4.2 (A1–A4), $\mathrm{CSD}(Q,C_{0})=0$ if and only if $Q=C_{0}$.
Proof is given in Appendix A.6.
Theorem 6.2 (Metrization of weak convergence).
Under Assumption A3, the copula-Stein discrepancy $\mathrm{CSD}(\cdot,C_{0})$ metrizes weak convergence: for any sequence $(Q_{n})$ in $\mathcal{P}([0,1]^{d})$ and any $Q\in\mathcal{P}([0,1]^{d})$,
$\mathrm{CSD}(Q_{n},C_{0})\to\mathrm{CSD}(Q,C_{0})\quad\Longleftrightarrow\quad Q_{n}\Rightarrow Q.$ (30)
Proof is provided in Appendix A.7.
6.2 Finite‐Sample Convergence and Optimality (Detailed Proof)
Theorem 6.3 (Finite‐Sample Convergence Rate).
Under Assumptions 4.2, one has
and hence $\mathbb{E}[\widehat{\mathrm{CSD}}_{n}^{2}]=O(n^{-1})$.
Full proof is given in Appendix A.8.
Corollary 6.4 (Optimal High‐Probability Convergence Rate).
Under the unified assumptions, the CSD estimator satisfies
$\widehat{\mathrm{CSD}}_{n}=O_{P}\bigl(n^{-1/2}\bigr).$ (33)
Proof is provided in Appendix A.9.
6.3 Sensitivity to Tail Dependence
Theorem 6.5 (Tail-dependence sensitivity).
$\mathrm{CSD}_{C_{2}}(C_{1})\;>\;0.$ (34)
Detailed proof is given in Appendix A.10.
Theorem 6.6 (Minimax risk for the squared discrepancy).
Let $\widehat{\mathrm{CSD}}_{n}^{2}$ be any estimator of $D^{2}(Q)\equiv\mathrm{CSD}(Q,C_{0})^{2}$ based on an i.i.d. sample of size $n$ from $Q$. Work over the local total-variation neighborhood
$\mathcal{Q}\equiv\Big\{Q:d_{\mathrm{TV}}(Q,C_{0})\leq\kappa n^{-1/2}\Big\}.$ (35)
(For the lower bound we do not restrict $\mathcal{Q}$ to copulas; the two-point construction below produces nearby alternatives dominated by $C_{0}$.) Under A1A3 there exist constants $0<c\leq C<\infty$ such that
$c\,n^{-2}\ \leq\ \inf_{\widehat{\mathrm{CSD}}_{n}^{2}}\ \sup_{Q\in\mathcal{Q}}\ \mathbb{E}_{Q}\big[\big(\widehat{\mathrm{CSD}}_{n}^{2}-D^{2}(Q)\big)^{2}\big]\ \leq\ C\,n^{-2}.$ (36)
Detailed proof is provided in Appendix A.11.
6.4 Summary of Theoretical Guarantees
  1. 1.
    Is a valid metric, as it metrizes weak convergence on the space of copula distributions (Theorem 6.2).
  2. 2.
    Is statistically optimal for finite samples, with its empirical estimator converging at the minimax optimal rate of $O_{P}(n^{-1/2})$ (Theorem 6.3, Corollary 6.4, and Theorem 6.6).
  3. 3.
    Is sensitive to fine-grained dependence, provably distinguishing between copulas with different tail dependence coefficients (Theorem 6.5).
Remark 6.7 (Extension).
Although our exposition focuses on Archimedean copulas for concreteness, the CSD framework relies only on the unified regularity conditions of the copula score $s(u)=\nabla\log c(u)$, and therefore extends to arbitrary copulas with sufficiently regular densities. We provide the general Copula-Stein operator and kernel formula, and show how the main guarantees (zero-mean, metrization, and finite-sample rate) carry over, along with examples for Gaussian, vine, and mixture copulas, in Appendix B.
Remark 6.8 (Computational Complexity ).
CSD is computationally tractable: with a separable $C^{2}$ base kernel, each Copula–Stein kernel evaluation costs $O(d)$, giving an exact estimator computable in $O(n^{2}d)$ time, while a random-feature approximation reduces the cost to $O(nmd)$ (nearly linear in $n$ when $m\ll n$). Full derivations and implementation details, including the exact-kernel routine, random-feature algorithm, and GPU and streaming variants, are provided in Appendix C.
7 Numerical Experiments
7.1 Experiment 1: empirical level and power (bivariate)
We study the finite-sample behavior of the CSD goodness-of-fit test in dimension $d=2$. Given i.i.d. pseudo-observations $U_{1},\dots,U_{n}\in(0,1)^{2}$, we compute the V-statistic
$\widehat{\mathrm{CSD}}_{n}^{2}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}),$ (37)
where $k_{C_{0}}$ is the Copula-Stein kernel defined in Prop. 5.2. We use the boundary-conditioned product RBF kernel with $b(u)=\prod_{j=1}^{2}u_{j}(1-u_{j})$ and choose the bandwidth $h$ by the median heuristic.
[…]
Table 1: Experiment 1 (bivariate): empirical Type I error and power for the CSD test ($\alpha=0.05$, $R=500$, $B=1000$).
$n$ Type I (Gumbel $\theta_{g}=2$) Power (Clayton $\theta_{c}=2$) Power (Gumbel $\theta_{\text{alt}}=2.4$)
200 0.046 0.104 0.748
500 0.064 0.268 1.000
1000 0.064 0.426 1.000
Overall, from Table 1, we see that the test exhibits near-nominal Type I error at $\alpha=0.05$ across sample sizes. We observe a slight liberal tendency at larger $n$ (Type I $\approx 0.064$ at $n\in\{500,1000\}$), which we attribute to finite-sample and Monte Carlo variability from bootstrap calibration and numerical score evaluation.
Power increases with $n$ for both alternatives: the test strongly detects within-family misspecification of the Gumbel parameter (power $0.748$ at $n=200$ and $1.0$ by $n=500$), while power against the Clayton alternative with matched Kendall’s $\tau$ is more moderate, reflecting a more subtle departure from the null at fixed $\tau$.
7.2 Experiment 2: random-feature approximation accuracy for $\mathrm{CSD}^{2}$
We evaluate the accuracy of the random-feature approximation used for scalable CSD. We generate a bivariate sample of size $n=5000$ from a Gumbel copula with Kendall’s $\tau=0.5$ (so $\theta=2$), compute the exact $\widehat{\mathrm{CSD}}_{n}^{2}$ using the Copula-Stein kernel (Prop. 5.6), and compare it to the random-feature approximation based on Algorithm 3. The bandwidth $h$ is set by the median heuristic. For each number of random features $m$, we repeat the random-feature construction $50$ times (independent draws of random features) and report the mean and standard deviation of the approximation and its relative error $|\mathrm{approx}-\mathrm{exact}|/|\mathrm{exact}|$.
Table 2: Experiment 2: random-feature approximation accuracy for $\mathrm{CSD}^{2}$ ($n=5000$, $\tau=0.5$, 50 independent random-feature draws per $m$). Exact $\widehat{\mathrm{CSD}}_{n}^{2}=5.893\times 10^{-4}$.
$m$ Approx. mean Approx. s.d. Rel. err. mean Rel. err. s.d.
250 $5.928\times 10^{-4}$ $2.60\times 10^{-5}$ $3.53\times 10^{-2}$ $2.66\times 10^{-2}$
500 $5.835\times 10^{-4}$ $2.12\times 10^{-5}$ $94\times 10^{-2} 2.2$$26\times 10^{-2}$
1000 $5.890\times 10^{-4}$ $1.51\times 10^{-5}$ $2.05\times 10^{-2}$ $1.50\times 10^{-2}$
2000 $5.894\times 10^{-4}$ $1.02\times 10^{-5}$ $1.34\times 10^{-2}$ $1.08\times 10^{-2}$
5000 $5.887\times 10^{-4}$ $6.72\times 10^{-6}$ $9.76\times 10^{-3}$ $5.83\times 10^{-3}$
10000 $5.899\times 10^{-4}$ $4.23\times 10^{-6}$ $5.90\times 10^{-3}$ $4.12\times 10^{-3}$
Figure 1: Experiment 2: mean relative error of the random-feature approximation versus the exact $\widehat{\mathrm{CSD}}_{n}^{2}$, as a function of the number of random features $m$ (log-log scale).
From Table 2 and Figure 1, we see that as $m$ increases, the approximation rapidly concentrates around the exact value: the mean relative error decreases from about $3.5\%$ at $m=250$ to below $1\%$ by $m=5000$, and to approximately $0.6\%$ at $m=10000$, supporting the accuracy of the scalable random-feature implementation.
8 Conclusion and Future Work
In this work, we introduced the Copula-Stein Discrepancy (CSD), a dependence-structure-aware framework for copula goodness-of-fit testing built by constructing a Stein operator directly on copula densities. Our theory provides a comprehensive treatment of Stein methods in the copula domain: CSD metrizes weak convergence on copula distributions, achieves the minimax-optimal convergence rate of $O_{P}(n^{-1/2})$, and is provably sensitive to tail-dependence differences via explicit lower bounds. For Archimedean copulas, we derived closed-form expressions for the CSD kernel, and our framework applies more broadly to copulas satisfying mild regularity conditions (including elliptical, vine, and mixture constructions).
We also addressed computational bottlenecks via exact computation with linear scaling in dimension and random-feature approximations that reduce the $O(n^{2})$ cost to nearly linear time. Our numerical results support these claims: Experiment 1 shows near-nominal Type I error at $\alpha=0.05$ with increasing power as $n$ grows, and Experiment 2 demonstrates that the random-feature approximation rapidly concentrates around the exact $\widehat{\mathrm{CSD}}_{n}^{2}$ as the number of features increases.
The framework developed here suggests several directions for further research. Immediate extensions include (i) a CSD-based two-sample test for comparing dependence structures, (ii) using CSD as an objective for copula parameter estimation as a robust alternative to maximum likelihood, and (iii) discrepancy-based model selection across copula families. On the methodological side, promising directions include incorporating CSD as a regularizer in deep generative models (e.g., VAEs/GANs) to improve dependence fidelity, extending the theory to high-dimensional regimes where $d$ grows with $n$, and developing streaming/online variants for real-time dependence monitoring.
Finally, broader empirical validation remains important, including benchmark comparisons against established copula GOF tests, real-data studies in domains where tail dependence is central (e.g., finance, hydrology, neuroscience), and open-source implementations in R and Python to support adoption and reproducibility.
Impact Statement
CSD is a statistical tool for assessing the goodness-of-fit of copula models, with an explicit focus on whether a proposed model captures multivariate dependence, including tail dependence, rather than on marginal behavior. If adopted, it could help researchers and practitioners more reliably validate dependence models in areas where joint extremes matter (for example, finance or hydrology), and it includes computational strategies intended to make such validation practical at larger sample sizes. However, like any model-checking method, CSD can be misused if its results are treated as a substitute for domain judgment or if it is deployed in high-stakes settings without careful attention to data quality, uncertainty, and downstream consequences. In particular, a well-fitting dependence model does not imply a system is fair or appropriate for decision-making, and failures can arise from distribution shift or incorrect preprocessing rather than the copula family alone. We view CSD as a diagnostic component within a broader responsible modeling workflow, where conclusions are supported by sensitivity analyses, transparent reporting, and domain-informed validation.
Appendix A Proofs
In this section, we provide the proofs of the lemmas, propositions, theorems and corollaries stated in the main text.
A.1 Proof of Lemma 5.2
Proof.
For each $j$,
$\int\partial_{u_{j}}g_{j}(u)\,c_{0}(u)\,du=\Big[g_{j}(u)\,c_{0}(u)\Big]_{u_{j}=0}^{u_{j}=1}-\int g_{j}(u)\,\partial_{u_{j}}c_{0}(u)\,du.$ (38)
Boundary vanishing implies the bracketed term is 0, and $\partial_{u_{j}}c_{0}=s_{j}c_{0}$. Summing over $j$ yields $\int\mathcal{T}_{c_{0}}g\,c_{0}\,du=0$, i.e. $\mathbb{E}_{C_{0}}[\mathcal{T}_{c_{0}}g(U)]=0$. ∎
A.2 Proof of Proposition 5.3
Proof.
Let $\mathcal{H}$ be the scalar RKHS with kernel $k$, so $\mathcal{H}^{d}$ is the $d$-fold product with $\langle g,h\rangle_{\mathcal{H}^{d}}=\sum_{j=1}^{d}\langle g_{j},h_{j}\rangle_{\mathcal{H}}$. By reproducing and closure under partials, $\partial_{u_{j}}g_{j}(u)=\langle g_{j},\,\partial_{u_{j}}k(u,\cdot)\rangle_{\mathcal{H}}$ and $g_{j}(u)=\langle g_{j},\,k(u,\cdot)\rangle_{\mathcal{H}}$. Hence
$\mathcal{T}_{c_{0}}g(u)=\sum_{j=1}^{d}\Big\langle g_{j},\;\partial_{u_{j}}k(u,\cdot)+(\partial_{u_{j}}\log c_{0}(u))\,k(u,\cdot)\Big\rangle_{\mathcal{H}}=\big\langle g,\,\xi_{C_{0}}(u)\big\rangle_{\mathcal{H}^{d}},$ (39)
with component $\xi_{C_{0}}(u)_{j}:=\partial_{u_{j}}k(u,\cdot)+s_{j}(u)\,k(u,\cdot)\in\mathcal{H}$. Thus $\xi_{C_{0}}(u)\in\mathcal{H}^{d}$. ∎
A.3 Proof of Theorem 5.4
Proof.
By Lemma 5.2,
$\text{CSD}(C_{n},C_{0})=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\left|\frac{1}{n}\sum_{i=1}^{n}\mathcal{T}_{c_{0}}g(U_{i})\right|=\sup_{\|g\|\leq 1}\left|\left\langle g,\;\mu_{n}\right\rangle_{\mathcal{H}^{d}}\right|,\quad\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{C_{0}}(U_{i}).$ (40)
$\text{CSD}^{2}(C_{n},C_{0})=\|\mu_{n}\|_{\mathcal{H}^{d}}^{2}=\left\langle\frac{1}{n}\sum_{i}\xi_{C_{0}}(U_{i}),\;\frac{1}{n}\sum_{j}\xi_{C_{0}}(U_{j})\right\rangle_{\mathcal{H}^{d}}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\langle\xi_{C_{0}}(U_{i}),\xi_{C_{0}}(U_{j})\rangle_{\mathcal{H}^{d}},$ (41)
which identifies $k_{C_{0}}(u,v):=\langle\xi_{C_{0}}(u),\xi_{C_{0}}(v)\rangle_{\mathcal{H}^{d}}$. ∎
A.4 Proof of Proposition 5.6
Proof.
From the witness function definition in Proposition 5.3:
@@ Appendix A Proofs
A.5 Proof of Proposition 5.7
Proof.
The standard density form is
$c(u)\;=\;\psi^{(d)}(t)\;\prod_{k=1}^{d}\!\big(-\varphi^{\prime}(u_{k})\big),\qquad t=\sum_{k=1}^{d}\varphi(u_{k}).$ (59)
@@
$\frac{\partial}{\partial u_{j}}\log c(u)=\Big(\partial_{t}\log\psi^{(d)}(t)\Big)\,\frac{\partial t}{\partial u_{j}}\;+\;\frac{\partial}{\partial u_{j}}\log\!\big(-\varphi^{\prime}(u_{j})\big).$ (61)
@@
Since $\partial t/\partial u_{j}=\varphi^{\prime}(u_{j})$ and $\partial_{t}\log\psi^{(d)}(t)=\psi^{(d+1)}(t)/\psi^{(d)}(t)$ while $\frac{\partial}{\partial u_{j}}\log\!\big(-\varphi^{\prime}(u_{j})\big)=\varphi^{\prime\prime}(u_{j})/\varphi^{\prime}(u_{j})$, the claim follows. ∎
A.6 Proof of Theorem 6.1
Proof.
By Propositions 5.1–5.2, $\mathrm{CSD}(\cdot,\cdot)$ is exactly the maximum mean discrepancy (MMD) associated with the Copula-Stein kernel $k_{C_{0}}$:
$\mathrm{CSD}(Q,C_{0})=\mathrm{MMD}_{k_{C_{0}}}(Q,C_{0}).$ (62)
Under Assumption A3, $k_{C_{0}}$ extends to a bounded, continuous, $c$-universal kernel on $0,1]^{d}$. For such kernels on compact metric spaces, $\mathrm{MMD}_{k_{C_{0}}}$ is a metric on $\mathcal{P}([0,1]^{d})$, hence $\mathrm{CSD}(Q,C_{0})=0$ iff $Q=C_{0}$. ∎
A.7 Proof of Theorem 6.2
Proof.
By Propositions 5.1–5.2, $\mathrm{CSD}(Q,C_{0})=\mathrm{MMD}_{k_{C_{0}}}(Q,C_{0})$, where $k_{C_{0}}$ is bounded and continuous on $0,1]^{d}$ under A3. Therefore, for any $Q_{n},Q\in\mathcal{P}([0,1]^{d})$,
$\big|\mathrm{CSD}(Q_{n},C_{0})-\mathrm{CSD}(Q,C_{0})\big|=\big|\mathrm{MMD}_{k_{C_{0}}}(Q_{n},C_{0})-\mathrm{MMD}_{k_{C_{0}}}(Q,C_{0})\big|\leq\mathrm{MMD}_{k_{C_{0}}}(Q_{n},Q).$ (63)
Moreover, on a compact metric space, any bounded continuous kernel whose MMD separates probability measures metrizes weak convergence. Since $k_{C_{0}}$ is $c$-universal on $0,1]^{d}$ (Assumption A3), it is characteristic, and therefore $\mathrm{MMD}_{k_{C_{0}}}$ metrizes weak convergence on $\mathcal{P}([0,1]^{d})$ (Sriperumbudur et al., 2010; Simon-Gabriel et al., 2023). Hence,
$\mathrm{MMD}_{k_{C_{0}}}(Q_{n},Q)\to 0\quad\Longleftrightarrow\quad Q_{n}\Rightarrow Q,$ (64)
which proves the claim. ∎
A.8 Proof of Theorem 6.3
Proof.
We write
@@
together with Lemma 5.2, imply $\mathbb{E}[k(U_{i},U_{j})]=0$. Hence
@@
as claimed. ∎
A.9 Proof of Corollary 6.4
Proof.
Decompose $\widehat{\mathrm{CSD}}_{n}^{2}=D_{\text{diag}}+D_{\text{off}}$ with